OpenStudy (anonymous):

Suppose a parabola has vertex (5, –3) and also passes through the point (6, 1). Write the equation of the parabola in vertex form. a.) y = (x – 5)2 – 3 b.) y= 4(x – 5)2 – 3 c.) y = 4(x + 5)2 – 3 d.) y = 4(x – 5)2 + 3

OpenStudy (anonymous):

It's either a.) or c.).... But idk which one...

OpenStudy (anonymous):

Vertex form is y=a(x-h)^2+k h=5, k=-3, so y=a(x-5)^2-3 plug in x and y - (6,1) - to solve for a.

OpenStudy (anonymous):

(c) is definitely out because it has (x+5) instead of (x-5).

OpenStudy (anonymous):

So it'a a then? or b?

OpenStudy (anonymous):

Did you solve for a yet?

OpenStudy (anonymous):

No cuz I can't....

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

When you put in (6,1) you get \[\large 1=a(6-5)^2-3\]

OpenStudy (anonymous):

So is it a, or b......?

OpenStudy (anonymous):

If you are not able to solve the above linear equation, then you should not be attempting to work with quadratic equations.

OpenStudy (anonymous):

I can, I've just been working on this for so long and it's my last problem and I don't want to do it anymore....

OpenStudy (anonymous):

I understand. Just get up, walk around a little; drink some water, and come back to the problem refreshed. You'll see what to do.